OpenStudy (anonymous):

Evaluate without using a calculator: tan (330) My answer was (-sqrt 3)/3 Please check my answer...i'm awful at math :/ Thanks! :)

OpenStudy (chaise):

This is correct! :)

OpenStudy (anonymous):

Thanks! Um, but i don't understand my next problem which is csc (-405) Can you explain to me please?

OpenStudy (chaise):

I'm not too familiar with csc, but what I will tell you is that -405degrees is the same as -45 degrees, if that is hepful!

OpenStudy (anonymous):

csc (angle) = 1/ sin (angle) cosecant is just the reciprocal of the sine function.

OpenStudy (anonymous):

would the answer be \[\sqrt{-2}\]?

OpenStudy (anonymous):

sorry i meant -\[\sqrt{2}\]

OpenStudy (anonymous):

forgot the negative....

OpenStudy (chaise):

she did do the negative, it was just in the next line. she did the negative out of equation editor and the sqrt(2) in the equation editor which moves it down to the next line it's a little ambiguous, but it is there.

OpenStudy (anonymous):

... so it's right? negative square root two? lol xD

OpenStudy (anonymous):

yes... see? don't every say you're awful at math again!!! good work. :)